A heating rated 100W is used to boil off 0.5kg of boiling water. the time to boil the water is ? (specific latent heat of vaporization= 2.30 × 10^6J/kg


1
Expert's answer
2022-03-07T10:19:11-0500

"P=\frac Qt=\frac{Lm}t,"

"t=\frac{Lm}P=3.2^h."